OpenStudy (anonymous):

The spanish club had 8 members who were 6th graders, 12 seventh graders, and 10 eighth graders. what percent of the club were seventh graders

OpenStudy (anonymous):

8 + 12 + 10 = 30 12 out of 30 = 12/30 = 2/5 = 40%

OpenStudy (anonymous):

thank you so much legomyego!